# Graph Chirp Signal and Graph Fractional Vertex-Frequency Energy Distribution

### v1

@@ IV-A Vertex-frequency energy distribution
IV-B Fractional vertex-frequency energy distribution
The energy of a signal in classical SP $\mathbf{x}$ is
@@ IV-B Fractional vertex-frequency energy distribution
Thus, we can define the GFED to further enhance the flexibility of signal analysis on graphs.
@@
$=$ $\sum_{m=1}^{N}x(n)\thinspace\overline{x(m)}\thinspace\overline{u_{k}^{a}(n)}\thinspace u_{k}^{a}(m)$ (17)
@@
Example 1: Consider a sensor network graph presented in Fig. 1(a). The graph chirp frequency $a=0.7$ is selected for the graph chirp signals, with initial frequencies chosen as follows. The graph signal $\mathbf{x}_{1}$, is defined on vertices 1 to 34 using the graph chirp signal $\mathbf{u}_{10}^{a}$, and on vertices 35 to 64 using $\mathbf{u}_{54}^{a}$, along with an additional $\mathbf{u}_{33}^{a}$. The graph signal $\mathbf{x}_{2}$, is defined on vertices 1 to 24 using $\mathbf{u}_{22}^{a}$, on vertices 25 to 34 using $\mathbf{u}_{7}^{a}$, and on vertices 35 to 64 using $\mathbf{u}_{42}^{a}$, along with the additional $\mathbf{u}_{33}^{a}$. The GEDs and GFEDs, along with the marginal properties are illustrated in Fig. 5.
@@
Figure 5: Sensor network graph structure and GFEDs. (a) Graph structure. (b) The GED of $\mathbf{x}_{1}$. (c) The GFED of $\mathbf{x}_{1}$ ($a=0.7$). (d) The GED of $\mathbf{x}_{2}$. (e) The GFED of $\mathbf{x}_{2}$ ($a=0.7$).
@@
Example 2: Consider a community network graph presented in Fig. 2(a). The graph chirp frequency $a=0.6$ is selected for the graph chirp signals, with initial frequencies chosen as follows. The graph signal $\mathbf{x}_{3}$, is defined on vertices 1 to 27 using the graph chirp signal $\mathbf{u}_{8}^{a}$, and on vertices 28 to 64 using $\mathbf{u}_{37}^{a}$, along with an additional $\mathbf{u}_{29}^{a}$. The graph signal $\mathbf{x}_{4}$, is defined on vertices 1 to 10 using $\mathbf{u}_{6}^{a}$, on vertices 11 to 27 using $\mathbf{u}_{15}^{a}$, and on vertices 28 to 64 using $\mathbf{u}_{51}^{a}$, along with the additional $\mathbf{u}_{29}^{a}$. The GEDs and GFEDs, along with the marginal properties are illustrated in Fig. 6.
@@
Figure 6: Community network graph structure and GFEDs. (a) Graph structure. (b) The GED of $\mathbf{x}_{3}$. (c) The GFED of $\mathbf{x}_{3}$ ($a=0.6$). (d) The GED of $\mathbf{x}_{4}$. (e) The GFED of $\mathbf{x}_{4}$ ($a=0.6$).
@@
As demonstrated in Examples 1 and 2, the GEDs fail to adequately capture the characteristics of graph chirp signals. In contrast, the GFEDs provide superior performance, effectively revealing vertex-frequency information, accurately identifying initial frequency components, and achieving greater energy concentration.
@@ V General Fractional Graph Distribution
To further extend the GGD, next, we introduce the GFGD.
@@
Also, the vertex moment property $\sum_{n=1}^{N}\sum_{k=1}^{N}n^{m}G_{\mathbf{x}}^{a}(n,k)=\sum_{n=1}^{N}n^{m}\left|x(n)\right|^{2}$ holds for the same condition.
@@
Also, the frequency moment property $\sum_{n=1}^{N}\sum_{k=1}^{N}n^{m}G_{\mathbf{x}}^{a}(n,k)=\sum_{k=1}^{N}k^{m}\left|\widehat{x}_{a}(k)\right|^{2}$ holds for the same condition.
Additionally, the GFGD can be rewritten as a dual form of (21) in the vertex-vertex domain as
@@
where $\varphi(m,n,t)$ is the kernel function in this domain, which has the same mathematical form with the frequency-frequency domain kernel in essence. Furthermore, if $\varphi(m,n,t)=\delta(m-n)$, the vertex marginal property is satisfied. Similarly, if $\sum_{n=1}^{N}\varphi(m,n,t)=1$, the frequency marginal property is satisfied.
@@
$\phi(p,k,q)=\frac{1}{s(p,q)}{\rm e}^{-\gamma\frac{|\lambda_{k}-\lambda_{q}|}{|\lambda_{p}-\lambda_{q}|}}$ (25)
@@
For the signals used in Examples 1–4, Fig. 7 plots the GGD with Choi-Williams kernel (GGD-CW) and GFGD with Choi-Williams kernel (GFGD-CW) of these four signals. Specifically, Figs. 3(a)–3(d) display the GGD-CW of the four signals, while Figs. 3(e)–3(h) display the GFGD-CW of the same signals.
Figure 7: The GGD-CWs and GFGD-CWs of the four signals. (a) The GGD-CW of $\mathbf{x}_{1}$. (b) The GGD-CW of $\mathbf{x}_{2}$. (c) The GGD-CW of $\mathbf{x}_{3}$. (d) The GGD-CW of $\mathbf{x}_{4}$. (e) The GFGD-CW of $\mathbf{x}_{1}$ ($a=0.7$). (f) The GFGD-CW of $\mathbf{x}_{2}$ ($a=0.7$). (g) The GFGD-CW of $\mathbf{x}_{3}$ ($a=0.6$). (h) The GFGD-CW of $\mathbf{x}_{4}$ ($a=0.6$).
As can be seen from Fig. 7, the GGD-CWs help reduce interference in the GEDs for signals shown in Figs. 5 and 6. However, they fail to correctly display the corresponding frequency components, as many other frequency components interfere the analysis. In contrast, the GFGD-CWs not only perform better in suppressing cross-terms in the GFEDs, but also provide a clearer representation of the vertex-fractional-frequency components, offering a more accurate depiction of the signals.
VI Detection of Graph Signals through Fractional vertex-frequency energy Distribution domain Filtering
Filtering and detecting play crucial roles in GSP, as they help in extracting relevant information while reducing unwanted noise. Typically, filtering is performed either in the vertex domain or in the frequency domain. In this section, we focus on signal detection through filtering techniques in the vertex-frequency domain, a hybrid approach that combines the strengths of both the vertex and frequency domains.
For the pure and noisy signals $\mathbf{x}$ and $\mathbf{y}$, we use the signal observation model $\mathbf{y}=\mathbf{x}+\mathbf{w}$, where $\mathbf{w}$ represents the additive noise. According to Wiener filter principle, a natural criterion to characterize the estimation accuracy is the MSE criterion
@@ VI Detection of Graph Signals through Fractional vertex-frequency energy Distribution domain Filtering
Similar to the convolution theorem in classical Fourier analysis, in GSP, we can utilize the convolution operator $\star$ to model the filtering process. Specifically, the estimated GFED $E_{\widetilde{\mathbf{x}}}^{a}$ is obtained through the convolution of the GFED of the noisy signal $\mathbf{y}$ and the filter $\mathbf{H}$ in vertex-frequency domain as
@@
If the original signal $\mathbf{x}$ is deterministic, $\mathbf{w}$ is zero-mean, complex circular Gaussian noise with variance $\sigma^{2}$, and $\mathbf{x}$ and $\mathbf{w}$ are independent, the filter transfer function matrix in the frequency-frequency domain can be derived to minimize the MSE as
$\widehat{\mathbf{H}}=\left(\mathbb{E}\left\{\widehat{E^{a}_{\mathbf{y}}}\circ\overline{\widehat{E^{a}_{\mathbf{y}}}}\right\}\right)^{-\circ}\circ\left(\widehat{E_{\mathbf{x}}}\circ\mathbb{E}\left\{\overline{\widehat{E^{a}_{\mathbf{y}}}}\right\}\right).$ (31)
Refer to Appendix A for the detailed derivation of (31).
@@
Thus, the optimal filter in the vertex-frequency domain can be expressed as (32).
@@
$H(n,k)=\sum_{l}\frac{\left|\widehat{E^{a}_{\mathbf{x}}}(l,k)\right|^{2}+\sigma^{2}\widehat{E^{a}_{\mathbf{x}}}(l,k)\sum_{i=1}^{N}\left|U_{a}(i,k)\right|^{2}\overline{U(i,l)}}{\begin{aligned} &\Bigg{[}\left|\widehat{E^{a}_{\mathbf{x}}}(l,k)\right|^{2}+\sigma^{2}\sum_{i}\sum_{j}x(i)\overline{\widehat{x}_{a}(k)}\overline{U_{a}(i,k)}\left|U_{a}(j,k)\right|^{2}\overline{U(i,l)}U(j,l)+\sigma^{2}\left|\sum_{i}x(i)\overline{U_{a}(i,k)}\overline{U(i,l)}\right|^{2}\\ &+\sigma^{2}\sum_{i}\left|\widehat{x}_{a}(k)\right|^{2}\left|U_{a}(i,k)\right|^{2}\left|U(i,l)\right|^{2}+\sigma^{2}\sum_{i}\sum_{j}\overline{x(j)}\widehat{x}_{a}(k)\left|U_{a}(i,k)\right|^{2}U_{a}(j,k)\overline{U(i,l)}U(j,l)\\ &+\sigma^{4}\left|\sum_{i}\left|U_{a}(i,k)\right|^{2}U(i,l)\right|^{2}+\sigma^{4}\sum_{i}\left|U_{a}(i,k)\right|^{2}\left|U(i,l)\right|^{2}+2\sigma^{4}\sum_{i}\left|U_{a}(i,k)\right|^{4}\left|U(i,l)\right|^{2}\Bigg{]}\end{aligned}}\thinspace U(n,l)$ (32)
Refer to Appendix B for the detailed derivation of (32).
@@
By employing the optimal filter, the minimum MSE is attained in the vertex-fractional-frequency domain, resulting in optimal detection and filtering performance. The subsequent section presents experimental results that validate this optimal performance.

### v2

@@ IV-A Vertex–frequency energy distribution
IV-B Fractional vertex–frequency energy distribution
The energy of a signal $\mathbf{x}$ in classical SP is
@@ IV-B Fractional vertex–frequency energy distribution
Thus, we can define the GFED to further enhance the flexibility of signal analysis on graphs. In particular, since the GFT operator used in the GED is not well–suited for handling graph chirp signals, we propose replacing it with the GFRFT operator, which is more appropriate for analyzing such signals. This substitution allows the GFED to better capture the joint vertex–fractional-frequency characteristics of graph chirps, thereby further enhancing the flexibility and effectiveness of signal analysis on graphs.
@@
$=$ $\sum_{m=1}^{N}x(n)\thinspace\overline{x(m)}\thinspace\overline{u_{k}^{a}(n)}\thinspace u_{k}^{a}(m).$ (27)
@@
Theorem 4 (Vertex–fractional-frequency domain energy concentration): Let $\mathbf{u}_{k_{0}}^{a_{0}}$ be a graph chirp signal with graph chirp rate $a_{0}$ and graph initial frequency $k_{0}$. Then its GFED with order $a$ satisfies
$E_{\mathbf{u}_{k_{0}}^{a_{0}}}^{a}(n,k)$ $=u_{k_{0}}^{a_{0}}(n)\,\overline{\widehat{\left(u_{k_{0}}^{a_{0}}\right)}_{a}(k)}\,\overline{u_{k}^{a}(n)}$
$=u_{k_{0}}^{a_{0}}(n)\,\overline{u_{k_{0}}^{a_{0}-a}(k)}\,\overline{u_{k}^{a}(n)}.$ (30)
In particular, when the graph chirp rate is equal to the fractional order, i,e, $a_{0}=a$, we have
$E_{\mathbf{u}_{k_{0}}^{a}}^{a}(n,k)=\begin{cases}\left|\mathbf{u}_{k_{0}}^{a}(n)\right|^{2}&k=k_{0}\\ 0&k\neq k_{0}\end{cases}.$ (31)
Proof: See Appendix A.
Notably, while GED fail to exhibit energy concentration for graph chirp signals, the GFED achieves perfect localization when the chirp rate $a_{0}$ matches the GFED order $a$, all energy is concentrated at the fractional frequency $k_{0}$. This result corresponds to the Property 4 of chirp signals. However, as explicitly pointed out by Stankovic and Sejdic in [24], the WD is not suitable for the graph framework extension. Therefore, in the graph setting, we consider the GFED as a functional analogue of the fractional WD/RD.
Although in classical SP the FRFT is widely interpreted as a rotation operator in the time–-frequency plane, effectively rotating the WD. However, this geometric interpretation can not extend to GSP, which operates in a non-Euclidean domain where concepts such as rotation are not naturally defined. The reason the FRFT can be interpreted as a rotation operator stems from the fact that the FT matrix $\mathbf{F}$ is a power–idempotent operator, satisfying $\mathbf{F}^{4}=\mathbf{I}_{N}$. In contrast, the GFT matrix is generally not power-idempotent, due to the irregularity of the underlying graph structure. If there exists an integer $a$ such that the GFT matrix satisfies $(\mathbf{F}_{G})^{a}=\mathbf{I}_{N}$, the integer $a$ can be interpreted as a form of generalized periodicity. However, such cases are rare. Therefore, the GFRFT does not produce a rotational effect analogous to that in the Euclidean setting, and the notion of rotation in the vertex–-frequency domain is not applicable in the framework of GSP.
@@
Example 1: Consider a sensor network graph presented in Fig. 4LABEL:sub@4a. The graph chirp frequency $a=0.7$ is selected for the graph chirp signals, with initial frequencies chosen as follows. The graph signal $\mathbf{x}_{1}$, is defined on vertices 1 to 24 using $\mathbf{u}_{22}^{a}$, on vertices 25 to 34 using $\mathbf{u}_{7}^{a}$, and on vertices 35 to 64 using $\mathbf{u}_{42}^{a}$, along with the additional $\mathbf{u}_{33}^{a}$. The GED and GFED, along with the marginal properties are illustrated in Fig. 4.
@@
Figure 4: Sensor network graph structure and GFEDs. (a) Graph structure. (b) The GED of $\mathbf{x}_{1}$. (c) The GFED of $\mathbf{x}_{1}$ ($a=0.7$).
@@
Example 2: Consider a community network graph presented in Fig. 5LABEL:sub@5a. The graph chirp frequency $a=0.6$ is selected for the graph chirp signals, with initial frequencies chosen as follows. The graph signal $\mathbf{x}_{2}$, is defined on vertices 1 to 27 using the graph chirp signal $\mathbf{u}_{8}^{a}$, and on vertices 28 to 64 using $\mathbf{u}_{37}^{a}$, along with an additional $\mathbf{u}_{29}^{a}$. The GED and GFED, along with the marginal properties are illustrated in Fig. 5.
@@
Figure 5: Community network graph structure and GFEDs. (a) Graph structure. (b) The GED of $\mathbf{x}_{2}$. (c) The GFED of $\mathbf{x}_{2}$ ($a=0.6$).
As observed in Figs. 4 and 5, the GFED representations offer clearer and more concentrated characterizations of graph chirp signals compared to the GED. Taking Example 2 as illustrated in Fig. 5, the difference between the two representations is evident. In Fig. 5LABEL:sub@5b, the GED of $\mathbf{x}_{2}$ displays several spread-out spectral components, with activations not only around the true initial frequencies (indices 8, 29, and 37), but also in unrelated regions such as around index 20, and between indices 40–60. These additional activations indicate spectral leakage and reduce interpretability. By contrast, as shown in Fig. 5LABEL:sub@5c, the GFED representation of $\mathbf{x}_{2}$ (with $a=0.6$) demonstrates a much more concentrated spectral profile, with energy sharply localized around the expected frequencies. This indicates that GFED provides a more accurate and compact vertex–fractional-frequency representation of graph chirp signals.
To move beyond purely visual comparison, we introduce Shannon entropy to quantify the concentration of energy. Given a normalized vertex–frequency distribution $\mathbf{D}_{\mathbf{x}}=\left[D(n,k)\right]_{N\times N}$ such that $\sum_{n}\sum_{k}\left|D(n,k)\right|^{2}=1$, the Shannon entropy is defined as [43]
$SE\left(\mathbf{D_{x}}\right)=-\sum_{n}\sum_{k}\left|D(n,k)\right|\mathrm{log}_{2}\left|D(n,k)\right|.$ (32)
A lower entropy value implies that the energy is more concentrated, while a higher entropy indicates a more dispersed energy distribution. Table I shows the entropy values for both GEDs and GFEDs of $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$. As presented in the Table, GFED consistently yields lower entropy values.
@@
It has been demonstrated in Examples 1 and 2 that the GEDs fail to adequately capture the characteristics of graph chirp signals. In contrast, the GFEDs provide superior performance, effectively revealing vertex–fractional-frequency information, accurately identifying initial frequency components, and achieving greater energy concentration.
TABLE I: Shannon entropy values of GEDs and GFEDs.
Signal Distribution Shannon entropy
$\mathbf{x}_{1}$ GED 55.864655.864655.8646
GFED ($a=0.7$) 54.5265
$\mathbf{x}_{2}$ GED 64.467164.467164.4671
GFED ($a=0.6$) 30.6718
@@ V General Fractional Graph Distribution
However, the GGD is also not well–suited for processing graph chirp signals. To address this limitation, we introduce the GFGD.
@@
Also, the vertex moment property
$\sum_{n=1}^{N}\sum_{k=1}^{N}n^{m}G_{\mathbf{x}}^{a}(n,k)=\sum_{n=1}^{N}n^{m}\left|x(n)\right|^{2}$ (36)
holds for the same condition.
@@
Also, the frequency moment property
$\sum_{n=1}^{N}\sum_{k=1}^{N}n^{m}G_{\mathbf{x}}^{a}(n,k)=\sum_{k=1}^{N}k^{m}\left|\widehat{x}_{a}(k)\right|^{2}$ (38)
holds for the same condition.
Additionally, the GFGD can be rewritten as a dual form of (34), expressed in the vertex–vertex domain as
@@
where $\varphi(m,n,t)$ is the kernel function in this domain, which has the same mathematical form with the frequency–frequency domain kernel in essence. Furthermore, if $\varphi(m,n,t)=\delta(m-n)$, the vertex marginal property is satisfied. Similarly, if $\sum_{n=1}^{N}\varphi(m,n,t)=1$, the frequency marginal property is satisfied.
@@
$\phi(p,k,q)=\frac{1}{s(p,q)}{\rm e}^{-\gamma\frac{|\lambda_{k}-\lambda_{q}|}{|\lambda_{p}-\lambda_{q}|}},$ (40)
@@
For the signals used in Examples 1–2, Fig. 6 plots the GGD with Choi-Williams kernel (GGD-CW) and GFGD with Choi-Williams kernel (GFGD-CW). As observed from Figs. 6LABEL:sub@6a and 6LABEL:sub@6c, the GGD-CW helps reduce interference in the GED. However, it still fails to accurately capture the true frequency components, as spectral artifacts remain and interfere with the interpretation. In contrast, the GFGD-CW achieves superior suppression of interference in the GFEDs and provides much clearer representations of the vertex–fractional-frequency features, offering a more accurate depiction of the signals. In addition, Table II shows the corresponding Shannon entropy values. The GFGD-CW consistently yields lower entropy values than the GGD-CW, indicating higher energy concentration and a more compact representation. These quantitative results are consistent with the visual analysis and confirm the effectiveness of GFGD-CW in reducing interference in the vertex–fractional-frequency domains.
Figure 6: The GGD-CWs and GFGD-CWs of the two signals. (a) The GGD-CW of $\mathbf{x}_{1}$. (b) The GFGD-CW of $\mathbf{x}_{1}$. (c) The GGD-CW of $\mathbf{x}_{2}$. (d) The GFGD-CW of $\mathbf{x}_{2}$.
TABLE II: Shannon entropy values of GGD-CWs and GFGD-CWs.
Signal Distribution Shannon entropy
$\mathbf{x}_{1}$ GGD-CW 55.500655.500655.5006
GFGD-CW ($a=0.7$) 48.2836
$\mathbf{x}_{2}$ GGD-CW 57.282457.282457.2824
GFGD-CW ($a=0.6$) 31.2431
VI Detection of Graph Signals through Fractional vertex–frequency energy Distribution domain Filtering
Filtering and detecting play crucial roles in GSP, as they help in extracting relevant information while reducing unwanted noise. Typically, filtering is performed either in the vertex domain or in the frequency domain. In this section, we focus on signal detection through filtering techniques in the vertex–fractional-frequency domain, a hybrid approach that combines the strengths of both the vertex and frequency domains.
For the pure and noisy signals $\mathbf{x}$ and $\mathbf{y}$, we use the signal observation model $\mathbf{y}=\mathbf{x}+\mathbf{w}$, where $\mathbf{w}$ represents the additive Gaussian noise. According to Wiener filter principle, a natural criterion to characterize the estimation accuracy is the MSE criterion
@@ VI Detection of Graph Signals through Fractional vertex–frequency energy Distribution domain Filtering
Similar to the convolution theorem in classical Fourier analysis, in GSP, we can utilize the convolution operator $\star$ to model the filtering process. Specifically, the estimated GFED $E_{\widetilde{\mathbf{x}}}^{a}$ is obtained through the convolution of the GFED of the noisy signal $\mathbf{y}$ and the filter $\mathbf{H}$ in vertex–fractional-frequency domain as
@@
Theorem 5: For the MSE minimization problem in (41), if the original signal $\mathbf{x}\in\mathbb{C}^{N}$ is deterministic, the noise $\mathbf{w}$ is zero-mean, circular complex Gaussian with variance $\sigma^{2}$, and $\mathbf{x}$ and $\mathbf{w}$ are statistically independent, then the filter transfer function matrix in the frequency–fractional-frequency domain that minimizes the MSE as
$\widehat{\mathbf{H}}=\left(\mathbb{E}\left\{\widehat{E^{a}_{\mathbf{y}}}\circ\overline{\widehat{E^{a}_{\mathbf{y}}}}\right\}\right)^{-\circ}\circ\left(\widehat{E_{\mathbf{x}}^{a}}\circ\mathbb{E}\left\{\overline{\widehat{E^{a}_{\mathbf{y}}}}\right\}\right).$ (46)
Proof: See Appendix B.
@@
Theorem 6: The optimal filter in the vertex–fractional-frequency domain for the MSE minimization problem in (41) can be expressed as (47).
@@
$H(n,k)=\sum_{l}\frac{\left[{\left|\widehat{E^{a}_{\mathbf{x}}}(l,k)\right|^{2}+\sigma^{2}\widehat{E^{a}_{\mathbf{x}}}(l,k)\sum_{i=1}^{N}\left|U_{a}(i,k)\right|^{2}\overline{U(i,l)}}\right]U(n,l)}{\begin{aligned} &\Bigg{[}\left|\widehat{E^{a}_{\mathbf{x}}}(l,k)\right|^{2}+\sigma^{2}\sum_{i}\sum_{j}x(i)\overline{\widehat{x}_{a}(k)}\overline{U_{a}(i,k)}\left|U_{a}(j,k)\right|^{2}\overline{U(i,l)}U(j,l)+\sigma^{2}\left|\sum_{i}x(i)\overline{U_{a}(i,k)}\overline{U(i,l)}\right|^{2}\\ &+\sigma^{2}\sum_{i}\left|\widehat{x}_{a}(k)\right|^{2}\left|U_{a}(i,k)\right|^{2}\left|U(i,l)\right|^{2}+\sigma^{2}\sum_{i}\sum_{j}\overline{x(j)}\widehat{x}_{a}(k)\left|U_{a}(i,k)\right|^{2}U_{a}(j,k)\overline{U(i,l)}U(j,l)\\ &+\sigma^{4}\left|\sum_{i}\left|U_{a}(i,k)\right|^{2}U(i,l)\right|^{2}+\sigma^{4}\sum_{i}\left|U_{a}(i,k)\right|^{2}\left|U(i,l)\right|^{2}+2\sigma^{4}\sum_{i}\left|U_{a}(i,k)\right|^{4}\left|U(i,l)\right|^{2}\Bigg{]}\end{aligned}}.$ (47)
Proof: See Appendix C.
@@
By employing the optimal filter, the minimum MSE is attained in the vertex–fractional-frequency domain, resulting in optimal detection and filtering performance. The subsequent section presents experimental results that validate this optimal performance.
Computational cost: The proposed GFED domain filtering method (GFED-F) involves several key operations whose computational complexity we now analyze. For an input graph signal $\mathbf{x}$ of length $N$, the computation of the GFRFT, the GFED, and the filter transfer function matrix in the vertex–-fractional-frequency domain, each require a computational complexity of $\mathcal{O}(N^{2})$. The computation of the eigendecomposition of the graph Laplacian $\mathbf{L}$, the GFT of GFED, and the estimated GFED through matrix multiplications each require a computational complexity of $\mathcal{O}(N^{3})$. Therefore, the overall computational complexity of the proposed algorithm is dominated by these cubic-time operations and is $\mathcal{O}(N^{3})$.
Our proposed GFED-F operates in the vertex–fractional-frequency domain based on the GFED. Specifically, the filtering operation is conducted in the frequency–fractional-frequency domain by projecting the GFED onto the GFT basis, thereby enabling a structured and interpretable filtering process. The filter transfer function matrix $\widehat{\mathbf{H}}$ resides in the frequency–fractional-frequency domain. Moreover, the GFED-F method avoids dependence on the clean signal by leveraging the prior GFED $E^{a}_{\mathbf{x}}$.
The graph Wiener filtering method in [45]: The ideal graph Wiener filtering method aims to minimize the MSE objective
$\mathbb{E}\left\{\left\|\mathbf{H}_{w}\mathbf{y}-\mathbf{x}\right\|_{2}^{2}\right\},$ (48)
where $\mathbf{H}_{w}$ is a spatial domain filter. Under the assumption that the observation covariance $\mathbb{E}\left\{\mathbf{y}\mathbf{y}^{\mathrm{H}}\right\}$ is invertible and that the original graph signal $\mathbf{x}$ is deterministic, the solution is given in closed form as
$\mathbf{H}_{w}=\frac{\mathbf{x}\mathbf{x}^{\mathrm{H}}}{\sigma^{2}+\|\mathbf{x}\|^{2}},$ (49)
which requires access to the pure signal $\mathbf{x}$. This is a non-parametric filtering approach performed entirely in the vertex domain with a computational complexity of $\mathcal{O}(N^{2})$.
The optimal GFRT domain filtering method (OGFRFT-F) [46]: The OGFRFT-F minimizes the MSE objective
$\mathbb{E}\left\{\left\|\mathbf{F}_{G}^{-a}\mathbf{H}_{o}\mathbf{F}_{G}^{a}\mathbf{y}-\mathbf{x}\right\|_{2}^{2}\right\},$ (50)
where $\mathbf{H}_{o}$ is a diagonal matrix in the GFRFT domain. This method performs filtering directly in the spectral domain defined by the GFRFT basis and is a parametric spectral domain filtering method. In essence, OGFRFT-F can be regarded as a constrained version of graph Wiener filtering, where the filter is restricted to be diagonal in the GFRFT domain. OGFRFT-F has a computational complexity of $\mathcal{O}(N^{4})$.
The GED-based Wiener filtering method (GED-WF) [47]: The GED-WF minimizes the MSE objective
$\mathbb{E}\left\{\left\|\left(\mathbf{F}_{G}^{-1}\circ\mathbf{H}_{g}\right)\mathbf{F}_{G}\mathbf{y}-\mathbf{x}\right\|_{2}^{2}\right\},$ (51)
where $\mathbf{H}_{g}$ is a filter defined in the vertex–frequency domain. (51) is mathematically equivalent to (48). However, the graph Wiener filtering method requires access to the clean signal $\mathbf{x}$, which limits its practical applicability. To address this limitation, some algorithms are proposed in [47] to estimate the filter $\mathbf{H}_{g}$ in the vertex–frequency domain without requiring knowledge of $\mathbf{x}$. These methods project the signal onto the GFT basis for filtering and maintain a computational complexity of $\mathcal{O}(N^{3})$.
In summary, both OGFRFT-F and GED-WF can be regarded as variants of the ideal graph Wiener filtering framework. In contrast, our proposed GFED-F method is fundamentally different in its formulation and mechanism, and thus also intrinsically distinct from OGFRFT-F and GED-WF. Table III summarizes the key differences among the four methods in terms of prior knowledge, projection basis, operational and filtering domains, and computational complexity.
TABLE III: Comparison of the four filtering methods.
Method Prior Knowledge Projection Basis Operational Domain Filtering Domain Computational Complexity
GFED-F $E^{a}_{\mathbf{x}}$, $\sigma$ GFT basis vertex–fractional-frequency domain frequency–fractional-frequency domain $\mathcal{O}(N^{3})$
graph Wiener filtering method $\mathbf{x}$, $\sigma$ canonical basis spatial domain spatial domain $\mathcal{O}(N^{2})$
OGFRFT-F $\mathbf{xx}^{\rm H}$, $\sigma$ GFRFT basis spatial domain fractional-frequency domain $\mathcal{O}(N^{4})$
GED-WF prior-free GFT basis spatial domain vertex–frequency domain $\mathcal{O}(N^{3})$
    $\mathbf{x}$ is deterministic.